# Spectrum-Aware Debiasing: A Modern Inference Framework with Applications to Principal Components Regression

### v4

@@ 1.1 Intuition via ridge estimator
Recall that we seek a debiased estimator of the form $\hat{\bm{\beta}}^{u}=\hat{\bm{\beta}}+\widehat{\mathsf{adj}}^{-1}\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})$. Suppose we plug in (5), leaving $\widehat{\mathsf{adj}}$ unspecified for the moment. If we denote the singular value decomposition of $\mathbf{X}$ to be $\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}$, we obtain that
@@
We operate under this assumption since it ensures ()($\star$)( ⋆ ) holds and our Spectrum-Aware adjustment turns out to be the correct debiasing strategy in this setting. Meanwhile, the degrees-of-freedom adjustment [12] yields the correction factor
@@
1.2 Right-rotationally invariant designs
Roughly speaking, Assumption 1 lands $\mathbf{X}$ in the class of right-rotationally invariant designs (Definition 2.1). Our arguments in the preceding section indicate that right-rotational invariance is a more fundamental assumption for debiasing than prior Gaussian/sub-Gaussian assumptions in the literature. Since it preserves the spectral information of $\mathbf{X}^{\top}\mathbf{X}$, we expect methods developed under this assumption to exhibit improved robustness when applied to designs that may not satisfy the right-rotational invariance property or designs observed in real data. We demonstrate this via Figure 3, bottom-left, where we conduct experiments with our PCR-Spectrum-Aware Debiasing on designs arising from six real datasets spanning image data, financial data, socio-economic data, and so forth.
Indeed varied research communities realized the strength of such designs, as demonstrated by the recent surge in literature in this space [85, 92, 86, 26, 27, 74, 108, 77, 6, 71, 72, 64, 73, 88, 62, 97, 34, 59, 76]. In particular, [27, 96] established that properties of high-dimensional systems proven under such designs continue to hold for a broad class of designs (including nearly deterministic designs as observed in compressed sensing [23]) as long as they satisfy certain spectral properties. In fact, the universality class for such designs is far broader than that for Gaussians, suggesting that these may serve as a solid prototype for modeling high-dimensional phenomena arising in non-Gaussian data. Despite such exciting developments, there are hardly any results when it comes to debiasing or inference under such designs (with the exception of [84] which we discuss later). This paper develops this important theory and methodology.
Despite the generality of right-rotationally invariant designs, studying these presents new challenges. For starters, analogs of the leave-one-out approach [67, 66, 9, 29, 104, 28, 82, 81, 21, 56] and Stein’s method [79, 20, 12, 10, 11, 2], both of which form fundamental proof techniques for Gaussian designs, are nonexistent or under-developed for this more general class. To mitigate this issue, we resort to an algorithmic proof strategy that the senior authors’ earlier work and that of others have used in the context of Gaussian designs. To study $\hat{\bm{\beta}}^{u}$, we observe that it depends on the regularized estimator $\bm{\hat{\beta}}$. However, $\bm{\hat{\beta}}$ does not admit a closed form in general, thus studying these turns out difficult. To circumvent this, we introduce a surrogate estimator that approximates $\bm{\hat{\beta}}$ in a suitable high-dimensional sense. We establish refined properties of these surrogates and use their closeness to our estimator of interest to infer properties about the latter. Prior literature has invoked this algorithmic proof strategy for Gaussian designs under the name of approximate message passing theory [24, 7, 82, 81, 107]. In case of right-rotationally invariant designs, we create the surrogate estimators using vector approximate message passing (VAMP) algorithms [74] (see details in Section A.6 and B.1 from Appendix). However, unlike the Gaussian case, proving that these surrogates approximate $\bm{\hat{\beta}}$ presents deep challenges. We overcome this via developing novel properties of VAMP algorithms that can be of independent interest to the signal processing [85], probability [92], statistical physics [86], information and coding theory [74, 73, 97] communities that seek to study models with right-rotationally invariant designs in the context of myriad other problems.
Among the literature related to right-rotationally invariant designs, two prior works are the most relevant for us. Of these, [41] initiated a study of the risk of $\bm{\hat{\beta}}$ under right-rotationally invariant designs using the VAMP machinery. However, their characterization is partially heuristic, meaning that they assume certain critical exchange of limits is allowed and that limits of certain fundamental quantities exist. The former assumption may often not hold, and the latter is unverifiable without proof (see Remark 5.1 for further details). As a by-product of our work on debiasing, we provide a complete rigorous characterization of the risk of regularized estimators under right-rotationally invariant designs (Theorem 5.1) without these unverifiable assumptions. The second relevant work is [84], which conjectures a population version of a debiasing formula for the Lasso using non-rigorous statistical physics tools. To be specific, they conjecture a debiasing formula that involves unknown parameters related to the underlying limiting spectral distribution of the sample covariance matrix. This formula does not provide an estimator that can be calculated from the observed data. In contrast, we develop a complete data-driven pipeline for debiasing and develop a consistent estimator for its asymptotic variance.
@@ 1.3 Practical issues and PCR-Spectrum-Aware Debiasing
This results in an inflation bias since $\frac{d_{1}^{2}}{d_{1}^{2}+\lambda_{2}}>\frac{1}{p}\sum_{i=1}^{p}\frac{d_{i}^{2}}{d_{i}^{2}+\lambda_{2}}.$ Similar problems arise if $\bm{\beta}^{\star}$ aligns with other eigenvectors, and the resulting bias could lead to inflation or shrinkage depending on the set of aligned eigenvectors. We refer to this as the alignment issue.
Another common issue arises when the top few eigenvalues of the sample covariance matrix $\mathbf{X}^{\top}\mathbf{X}$ are significantly separated from the bulk of the spectrum. In this case, after centering the spectrum of $\mathbf{V}$, the variance of the “$\mathsf{unbiased\;component}$” in (10) will be large, making the debiasing procedure unstable. We refer to these eigenvalues as outlier eigenvalues.
@@ 1.4 Notable outcomes: Alignment Testing and Debiased PCR
Our theory for Spectrum-Aware Debiasing has two significant by-products. In modern high-dimensional inference, calculating the precise asymptotic risk of regularized estimators has emerged as a prominent research area. Technical tools such as approximate message passing [24, 8, 7, 104, 52, 81, 5, 36], the convex Gaussian min-max theorem [91, 80], and the cavity or leave-one-out method [67, 89, 29, 28, 9, 82, 81, 21, 56] have proven invaluable for this purpose. These tools have facilitated the discovery of novel high-dimensional phenomena that other mathematical techniques simply fail to capture [22, 29, 9, 28, 82, 81, 17, 107, 61, 60, 56, 109]. Consequently, they have inspired new high-dimensional estimators that outperform traditional ones by a margin [81, 19]. Despite such remarkable progress, these technical tools suffer a crucial limitation. They typically assume that the design matrices are random and independent of the true signal, implying that the PCs are random vectors in generic position relative to the true signal. To the best of our knowledge, a principled test to validate this assumption has so far eluded the literature. In this paper, we introduce the first formal hypothesis test for PC-signal alignment, utilizing our PCR-Spectrum-Aware approach (see Corollary 4.3 and the subsequent discussion). We hope this serves as a foundation for more systematic investigations into this issue, thereby enhancing the applicability of statistical methods developed based on approximate message passing algorithms/leave-one-out/convex Gaussian min-max theorems.
As a second outcome, our work contributes to an extensive and growing body of work on PCR methodologies [57, 49, 4, 47, 30, 1, 46, 106, 50, 78, 83, 15]. Similar to the ridge and Lasso estimators, the traditional PCR estimator exhibits shrinkage bias due to the discarding of low-variance PCs [35, 38, 39, 14, 25, 58]. To the best of our knowledge, no previous work has investigated how this bias can be eliminated in high dimensions and its implications for inference. We develop the first approach for debiasing the classical PCR estimator, complete with formal high-dimensional guarantees. We will next formally introduce Spectrum-Aware Debiasing and discuss its properties.
@@ 1.5 Organization
We organize the rest of the paper as follows. In Section 2, we introduce our assumptions and preliminaries. In Sections 3 and 4, we introduce our Spectrum-Aware and PCR-Spectrum-Aware methods with formal guarantees. In Section 5, we present our proof outline and technical novelties. Finally in Section 6, we conclude with potential directions for future work.
@@ 2.1 Design matrix, signal and noise
$\left|f\left(\mathbf{v}_{1},\ldots,\mathbf{v}_{k}\right)\right|\leq C\left(1+\left\|\left(\mathbf{v}_{1},\ldots,\mathbf{v}_{k}\right)\right\|^{2}\right)$ (11)
@@
where $\mathsf{D}^{2}$ has non-zero mean with compact support $\operatorname{supp}(\mathsf{D}^{2})\subseteq[0,\infty)$. We denote $d_{-}:=\min(x:x\in\operatorname{supp}(\mathsf{D}^{2}))$. Furthermore, we assume that as $p\to\infty$,
$d_{+}:=\lim_{p\to\infty}\max_{i\in[p]}d_{i}^{2}<+\infty.$ (13)
@@
Since our debiasing procedure relies on the spectrum of $\mathbf{X}^{\top}\mathbf{X}$, analyzing its properties requires a thorough understanding of the properties of $\mathsf{D}$ (from (12)), the limit of the empirical spectral distribution of $\mathbf{X}^{\top}\mathbf{X}$. Often these properties can be expressed using two important quantities—the Cauchy and the R-transform. We define these next. For technical reasons, we will define these transforms corresponding to the law of $-\mathsf{D}^{2}$.
@@
Under Assumption 2, let $G:(-d_{-},\infty)\to(0,\infty)$ and $R:(0,G(-d_{-}))\to(-\infty,0)$ be the Cauchy- and R-transforms of the law of $-\mathsf{D}^{2}$, defined as
@@ 2.2 Penalty function
We assume that $h:\mathbb{R}\mapsto[0,+\infty)$ is proper, closed and satisfies that for some $c_{0}\geq 0,\forall x,y\in\mathbb{R},t\in[0,1]$,
@@
Here, $c_{0}=0$ is equivalent to assuming $h$ is convex and $c_{0}>0$ to assuming $h$ is $c_{0}$-strongly convex. Furthermore, we assume that $h(x)$ is twice continuously differentiable except for a finite set $\mathfrak{D}$ of points, and that $h^{\prime\prime}(x)$ and $\operatorname{Prox}_{vh}^{\prime}(x)$ have been extended at their respective undefined points using Lemma 2.7 below.
@@
Fix any $v>0$. Under Assumption 4, $x\mapsto\operatorname{Prox}_{vh}(x)$ is continuously differentiable at all but a finite set $\mathcal{C}$ of points. Extending functions $x\mapsto h^{\prime\prime}(x)$ and $x\mapsto\operatorname{Prox}_{vh}^{\prime}(x)$ on $\mathfrak{D}$ and $\mathcal{C}$ by +$\infty$+ ∞ and 00 respectively, we have that for all $x\in\mathbb{R}$,
@@
After the extension, for any $w>0$, $x\mapsto\frac{1}{w+h^{\prime\prime}\left(\operatorname{Prox}_{vh}(x)\right)}$ is piecewise continuous with finitely many discontinuity points on which it takes value 00.
@@
The following assumption is analogous to [12, Assumption 3.1] for the Gaussian design: we require either $h$ to be strongly convex or $\mathbf{X}^{\top}\mathbf{X}$ to be non-singular with smallest eigenvalues bounded away from 00.
@@
Either $c_{0}>0$ or $\lim_{p\to\infty}\min_{i\in p}(d_{i}^{2})\geq c_{1}$ for some constant $c_{1}>0$.
@@ 2.3 Fixed-point equation
Recall our discussion from Section 1.2. We study the regularized estimator $\bm{\hat{\beta}}$ by introducing a more tractable surrogate $\bm{\hat{\beta}}^{t}$. Later we will see that we construct this surrogate using an iterative algorithmic scheme known as Vector Approximate Message Passing algorithm (VAMP) [74]. Thus to study the surrogate, one needs to study the VAMP algorithm carefully. One can describe the properties of this algorithm using a system of fixed point equations in four variables. We use $\gamma_{*},\eta_{*},\tau_{*},\tau_{**},\in(0,+\infty)$ to denote these variables, and define the system here:
$\frac{\gamma_{*}}{\eta_{*}}=\mathbb{E}\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\mathsf{B}^{\star}+\sqrt{\tau_{*}}\mathsf{Z}\right)$ (20a)
$\tau_{**}=\frac{\eta_{*}^{2}}{\left(\eta_{*}-\gamma_{*}\right)^{2}}\left[\mathbb{E}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\mathsf{B}^{\star}+\sqrt{\tau_{*}}\mathsf{Z}\right)-\mathsf{B}^{\star}\right)^{2}-\left(\frac{\gamma_{*}}{\eta_{*}}\right)^{2}\tau_{*}\right]$ (20b)
$\gamma_{*}=-R\left(\eta_{*}^{-1}\right)$ (20c)
@@
There exists a solution $\gamma_{*},\eta_{*},\tau_{*},\tau_{**}\in(0,+\infty)$ such that (20) holds.
@@
Under Assumptions 25, the existence of a solution implies uniqueness, as we show in Section A.5.2 from Appendix.
The following Proposition shows that Assumption 6 holds for any Elastic Net penalties (that are not Lasso). See proof in Section A.5.3 from Appendix.
Proposition 2.10 (Assumption 6 holds for Elastic Net).
Suppose that $\mathsf{D}^{2}$ and $\mathsf{B}^{\star}$ are as in Assumption 2 and 3. For the Elastic Net penalty $h$ defined in Example 2.8 with $\lambda_{2}>0$, Assumption 6 holds for any $\sigma^{2}>0$.
Remark 2.11 (Verifying Assumption 6).
Proving Assumption 6 is difficult for general penalties and $\mathsf{D}^{2}$. One therefore hopes to verify it on a case-by-case basis. For instance, when $\mathsf{D}^{2}$ follows the Marchenko-Pastur law and $h$ is the Lasso, [7, Proposition 1.3, 1.4] proved that (20) admits a unique solution. Similarly, [82] proved existence and uniqueness of the solution to an analogous fixed point equation in the context of logistic and probit regression. For general $\mathsf{D}^{2}$ in the context of linear regression, there are no rigorous results establishing that the system admits a solution. In Proposition 2.10, we take the first step and show for the first time that Assumption 6 indeed holds for the Elastic Net for a general $\mathsf{D}^{2}$ satisfying our assumptions. This result may be of independent interest: we anticipate that our proof can be adapted to other penalties beyond the Elastic Net; however, since our arguments require an explicit expression for the proximal operator, we limit our presentation to the Elastic Net. That said, one expects this to always hold true under Assumptions 25 (see e.g. [40, 41]).
@@
Assumption 7 serves as a technical condition to rule out degenerate scenarios where estimating $\sigma^{2}$ is impossible. For example, this condition is not satisfied when $n=p$ and $\mathbf{X}=\mathbf{I}_{p}$: in this case, our sole observation is $\mathbf{y}=\bm{\beta}^{\star}+\bm{\varepsilon}$ and it is indeed impossible to estimate $\sigma^{2}$. We provide a consistent estimator ifor the left-hand side of (21) in (39), facilitating the verification of Assumption 7.

### v6

@@ 1.1 Intuition via ridge estimator
[…]
@@
We operate under this assumption since it ensures $(\star)$ holds and our Spectrum-Aware adjustment turns out to be the correct debiasing strategy in this setting. Meanwhile, the degrees-of-freedom adjustment [13] yields the correction factor
@@ 1.2 Practical issues and PCR-Spectrum-Aware Debiasing
This results in an inflation bias since $\frac{d_{1}^{2}}{d_{1}^{2}+\lambda_{2}}>\frac{1}{p}\sum_{i=1}^{p}\frac{d_{i}^{2}}{d_{i}^{2}+\lambda_{2}}.$ Similar problems arise if $\bm{\beta}^{\star}$ aligns with other eigenvectors, and the resulting bias could lead to inflation or shrinkage depending on the set of aligned eigenvectors. We refer to this as the alignment issue. Another common issue arises when the top few eigenvalues of the sample covariance matrix $\mathbf{X}^{\top}\mathbf{X}$ are significantly separated from the bulk of the spectrum. In this case, after centering the spectrum of $\mathbf{V}$, the variance of the $\mathsf{unbiased\;component}$” in (10) will be large, making the debiasing procedure unstable. We refer to these eigenvalues as outlier eigenvalues.
@@ 1.3 Notable outcomes: Alignment Testing and Debiased PCR
Our theory for Spectrum-Aware Debiasing has two significant by-products. In modern high-dimensional inference, calculating the precise asymptotic risk of regularized estimators has emerged as a prominent research area. Technical tools such as approximate message passing [30, 9, 10, 122, 60, 97, 7, 43], the convex Gaussian min-max theorem [106, 95], random matrix theory [26, 27, 55, 25, 2, 70], and the cavity or leave-one-out method [80, 104, 36, 35, 11, 98, 97, 24, 64] have proven invaluable for this purpose. These tools have facilitated the discovery of novel high-dimensional phenomena that other mathematical techniques simply fail to capture [28, 36, 11, 35, 98, 97, 19, 124, 72, 71, 64, 125]. Consequently, they have inspired new high-dimensional estimators that outperform traditional ones by a margin [97, 21]. Despite such remarkable progress, these technical tools suffer a crucial limitation. They typically assume that the design matrices are random and independent of the true signal, implying that the PCs are random vectors in generic position relative to the true signal. To the best of our knowledge, a principled test to validate this assumption has so far eluded the literature. In this paper, we introduce the first formal hypothesis test for PC-signal alignment, utilizing our PCR-Spectrum-Aware approach (see Corollary 4.5 and the subsequent discussion). We hope this serves as a foundation for more systematic investigations into this issue, thereby enhancing the applicability of statistical methods developed based on approximate message passing algorithms/leave-one-out/convex Gaussian min-max theorems.
As a second outcome, our work contributes to an extensive and growing body of work on PCR methodologies [65, 58, 6, 56, 37, 3, 93, 16]. Similar to the ridge and Lasso estimators, the traditional PCR estimator exhibits shrinkage bias due to the discarding of low-variance PCs [42, 46, 47, 15, 31, 66]. To the best of our knowledge, no previous work has investigated how this bias can be eliminated in high dimensions and its implications for inference. We develop the first approach for debiasing the classical PCR estimator, complete with formal high-dimensional guarantees. We will next formally introduce Spectrum-Aware Debiasing and discuss its properties.
@@ 1.4 Organization
We organize the rest of the paper as follows. In Section 2, we introduce our assumptions and preliminaries. In Sections 3 and 4, we introduce our Spectrum-Aware and PCR-Spectrum-Aware methods with formal guarantees. Finally in Section 5, we conclude with potential directions for future work.
@@ 2.1 Design matrix, signal and noise
$\left|f\left(v_{1},\ldots,v_{k}\right)\right|\leq C\left(1+\left\|\left(v_{1},\ldots,v_{k}\right)\right\|^{2}\right)$ (11)
@@
where $\mathsf{D}^{2}$ has non-zero mean with compact support [^3] Throughout, we define support of a random variable $X$ as the smallest closed set $A$ such that $\mathbb{P}(X\in A)=1$. $\operatorname{supp}(\mathsf{D}^{2})\subseteq[0,\infty)$. We denote $d_{-}:=\min(x:x\in\operatorname{supp}(\mathsf{D}^{2}))$. Furthermore, we assume that as $p\to\infty$,
$d_{+}:=\limsup_{p\to\infty}\max_{i\in[p]}d_{i}^{2}<+\infty.$ (13)
@@
Since our debiasing procedure relies on the spectrum of $\mathbf{X}^{\top}\mathbf{X}$, analyzing its properties requires a thorough understanding of the properties of $\mathsf{D}$ (from (12)), the limit of the empirical spectral distribution of $\mathbf{X}^{\top}\mathbf{X}$. Often these properties can be expressed using two important quantities—the Cauchy and the R-transform. We define these next. For technical reasons, we will define these transforms corresponding to the law of $\mathsf{D}^{2}$.
@@
Under Assumption 1, let $G:(-d_{-},\infty)\to(0,\infty)$ and $R:(0,G(-d_{-}))\to(-\infty,0)$ be the Cauchy- and R-transforms of the law of $\mathsf{D}^{2}$, defined as
@@
The assumption on the signal $\bm{\beta}^{\star}\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{B}^{\star}$ may be relaxed using recent non-asymptotic AMP/VAMP theories [17, 68, 67]. We leave this to future works.
Remark 2.7.
@@ 2.2 Penalty function
We assume that $h:\mathbb{R}\mapsto[0,+\infty)$ is non-constant, proper and closed convex function. Furthermore, we assume that $h(x)$ is twice continuously differentiable except for a finite set $\mathfrak{D}$ of points, and that $h^{\prime\prime}(x)$ and $\operatorname{Prox}_{vh}^{\prime}(x)$ have been extended at their respective undefined points using Lemma 2.8 below.
Note that convexity of $h$ in Assumption 3 implies that for some $c_{0}\geq 0,\forall x,y\in\mathbb{R},t\in[0,1]$,
@@
Here, $h$ is said to be strongly convex if $c_{0}>0$.
@@
Fix any $v>0$. Under Assumption 3, $x\mapsto\operatorname{Prox}_{vh}(x)$ is continuously differentiable at all but a finite set $\mathcal{C}$ of points. Extending functions $x\mapsto h^{\prime\prime}(x)$ and $x\mapsto\operatorname{Prox}_{vh}^{\prime}(x)$ on $\mathfrak{D}$ and $\mathcal{C}$ by $\infty$ and 0 respectively, we have that for all $x\in\mathbb{R}$,
@@
After the extension, for any $w>0$, $x\mapsto\frac{1}{w+h^{\prime\prime}\left(\operatorname{Prox}_{vh}(x)\right)}$ is piecewise continuous with finitely many discontinuity points on which it takes value 0.
@@
The following assumption is analogous to [13, Assumption 3.1] for the Gaussian design: we require either $h$ to be strongly convex or $\mathbf{X}^{\top}\mathbf{X}$ to be non-singular with smallest eigenvalues bounded away from 0.
@@
Either $c_{0}>0$ or $d_{-}:=\lim_{p\to\infty}\min_{i\in p}(d_{i}^{2})\geq c_{1}$ for some constant $c_{1}>0$.
However, we show that for the Lasso with $h(x)=\lambda_{1}|x|,\lambda_{1}>0$ (where $c_{0}=0$), we may be able to drop the requirement that $d_{-}>0$ if the penalty strength $\lambda_{1}>0$ is sufficiently large. We defer the discussion to Section F.2.
@@ 2.3 Fixed-point equation
Our general approach to study the regularized estimator $\bm{\hat{\beta}}$ is by introducing a more tractable surrogate $\bm{\hat{\beta}}^{t}$. As detailed in the Appendix, Appendix C, we construct this surrogate using an iterative algorithmic scheme known as Vector Approximate Message Passing algorithm (VAMP) [88]. Thus to study the surrogate, one needs to study the VAMP algorithm carefully. One can describe the properties of this algorithm using a system of fixed point equations in four variables. We use $\gamma_{*},\eta_{*},\tau_{*},\tau_{**},\in(0,+\infty)$ to denote these variables, and define the system here:
$\frac{\gamma_{*}}{\eta_{*}}=\mathbb{E}\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}\left(\mathsf{B}^{\star}+\sqrt{\tau_{*}}\mathsf{Z}\right),$ (20a)
$\tau_{**}=\frac{\eta_{*}^{2}}{\left(\eta_{*}-\gamma_{*}\right)^{2}}\left[\mathbb{E}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\mathsf{B}^{\star}+\sqrt{\tau_{*}}\mathsf{Z}\right)-\mathsf{B}^{\star}\right)^{2}-\left(\frac{\gamma_{*}}{\eta_{*}}\right)^{2}\tau_{*}\right],$ (20b)
$\gamma_{*}=-R\left(\eta_{*}^{-1}\right),$ (20c)
@@
There exists a solution $\gamma_{*},\eta_{*},\tau_{*},\tau_{**}\in(0,+\infty)$ and $\eta_{*}>\gamma_{*}$ such that (20) holds.
We now provide sufficient conditions under which Assumption 5 holds. While the system of fixed-point equations (20) plays a central role in the theory of regularized high-dimensional linear regression with right-rotationally invariant designs [48, 49], there have been no rigorous mathematical results establishing the existence of solutions. The following proposition demonstrates that Assumption 5 holds for a class of strongly convex penalties satisfying Assumption 6.
Assumption 6.
We assume that the proximal operator $\operatorname{Prox}_{vh}(x)$ satisfies the following properties:
  • (i)
    Monotonicity. For any $b\in\mathbb{R},\alpha>0,v>0$,
    $\frac{\partial}{\partial v}\mathbb{E}\operatorname{Prox}_{vh}^{\prime}\left(b+\frac{v}{\alpha}\mathsf{Z}\right)\leq 0$ (21)
    where $\mathsf{Z}\sim N(0,1)$.
  • (ii)
    Asymptotic linearity. There exists continuous functions $r(v)$ defined on $v\in(0,+\infty)$ such that for each fixed $v>0$,
    $\lim_{x\rightarrow-\infty}\operatorname{Prox}^{\prime}_{vh}(x)=\lim_{x\rightarrow+\infty}\operatorname{Prox}^{\prime}_{vh}(x)=r(v)>0.$ (22)
Remark 2.10.
Ridge, Lasso, Elastic Net (cf.  Example 2.9) and Huber Norm penalty (cf. [121]), defined as, for tuning parameters $u,\delta>0,\lambda_{1}\geq 0,\lambda_{2}\geq 0$,
$h(x)=\lambda_{1}g(x)+\frac{\lambda_{2}}{2}x^{2},\quad g(x)=\begin{cases}\dfrac{u}{2}\,x^{2},&|x|\leq\delta,\\[6.0pt] u\delta\bigl(|x|-\tfrac{\delta}{2}\bigr),&|x|>\delta,\end{cases}$
all satisfy Assumption 6. See Proposition B.2.
The proof of Proposition 2.11 under strong convexity is in Section B.3 while the Lasso case is shown in Proposition H.3. For the Lasso case, Proposition 2.11 imposes a mild condition that $\mathsf{D}^{2}$ places nonzero mass at zero or it satisfies $d_{-}>0$. This rules out the edge case where $\mathsf{D}^{2}$ has no mass at zero but the eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$ are not bounded away from zero. When $\mathbf{X}$ has i.i.d. sub-Gaussian entries and $n/p\to\varsigma$, $\mathsf{D}^{2}$ follows the Marchenko–Pastur law; the condition then excludes the edge case $\varsigma=1$.
Proposition 2.11.
Let $\mathsf{D}^{2}$ be the random variable defined in Assumption 1, $h$ is strongly convex, satisfying Assumption 3 with $c_{0}>0$, and that proximal operator $\operatorname{Prox}_{vh}(x)$ satisfies Assumption 6. Then Assumption 5 holds.
Furthermore, under a mild condition on $\mathsf{D}^{2}$ — specifically, if it either has non-zero probability mass at zero or satisfies $d_{-}>0$ (cf. Assumption 13) — Assumption 5 also holds for the Lasso penalty, i.e., when $h$ is as in Example 2.9 with $\lambda_{1}>0$ and $\lambda_{2}=0$.
@@
Under Assumptions 14, the existence of a solution implies uniqueness, as we show in Section B.2 from Appendix.
@@
Assumption 7 serves as a technical condition to rule out degenerate scenarios where estimating $\sigma^{2}$ is impossible. For example, this condition is not satisfied when $n=p$ and $\mathbf{X}=\mathbf{I}_{p}$: in this case, our sole observation is $\mathbf{y}=\bm{\beta}^{\star}+\bm{\varepsilon}$ and it is indeed impossible to estimate $\sigma^{2}$. We provide a consistent estimator for the left-hand side of (23) in (67), facilitating the verification of Assumption 7.